OpenStudy (anonymous):

26,000 / .0013 solve in scientific notation form. I am getting 2.0 x 10^6, but it should be 2.0 x 10 ^7..... can someone explain ??

OpenStudy (tkhunny):

2.6 E4 1.3 E-3 2.6/1.3 = 2 E4/E-3 = E7 So, why are you getting E6?

OpenStudy (anonymous):

check where your put your decimal points.

OpenStudy (mathstudent55):

\( \dfrac{26000}{0.0013} \) \( = \dfrac{2.6 \times 10^4}{1.3 \times 10^{-3}} \) \(= 2 \times 10^{4 - (-3)} \) \(= 2 \times 10^7\)

OpenStudy (anonymous):

\[\dfrac{26,000}{0.0013}=\frac{260,000,000}{13}\]

OpenStudy (anonymous):

you get a 2 followed by 7 zeros

OpenStudy (anonymous):

Thx everyone!

OpenStudy (mathstudent55):

wlcm